OpenStudy (anonymous):

how do you solve x=3y-4 7=y-2x

OpenStudy (anonymous):

2 equal signs??

OpenStudy (anonymous):

What do x and y represent? It is a bit confusing for me.

OpenStudy (anonymous):

You can solve by a method called substitution

OpenStudy (p0sitr0n):

substitute x 7 = y-2(3y-4)

OpenStudy (anonymous):

x=3y-4 2x=y-7 x=.5(y-7) plug in one x into the other

OpenStudy (anonymous):

x is equal to 3y-4 therefore, in the second equation, write 7=y-2(3y-4)

OpenStudy (anonymous):

\[x = -17/5\] \[y = 1/5\]

OpenStudy (anonymous):

Solve this, then plug into the y value into first equaiton to find x

OpenStudy (anonymous):

Why is this place filled to the brim? o_0

OpenStudy (anonymous):

im so confused what?

OpenStudy (anonymous):

Some of you proponents of the substitution method might find this interesting:  It really shines when you have 3 or more variables.

OpenStudy (anonymous):

Try substituting the first equation into the second equation. x= >>> 3y-4 <<< 7=y-2x So: 7=y-2( >>> 3y-4 <<< ) 7=y-2(3y-4) See how I plugged in the value of x into the second equation?

OpenStudy (anonymous):

the lines intersect at \[-17/5 , 1/5\] though

OpenStudy (anonymous):

yea i already understand that its the whole solving part i dont get

OpenStudy (anonymous):

@alexarchival Show your work so far.

OpenStudy (anonymous):

i dont get any of this all i get is how to plug one x in for the other i dont know how to solve

OpenStudy (anonymous):

would it be 7=4y+6

OpenStudy (anonymous):

7 = y -2(3y-4) Distribute the -2: 7 = y -6y +8 7 = -5y + 8 -1 = -5y y = 1/5 Then plug that y into your original equation to get x.

OpenStudy (anonymous):

i got -3.4

OpenStudy (anonymous):

That's correct. :) Though, your teacher might want it in fraction form.